An ion engine has a 1T magnetic field and hydrogen plasma is to be short out at an E×B velocity of 1150km/s. how much internal electric field must be present in the plasma?
1
Expert's answer
2020-05-15T08:41:47-0400

The speed of electromagnetic field is given by the formula


"v=\frac{E}{B}"

where E is electric field intensity and B is magnetic field intensity.

"E=vB=(1)(1.15\cdot 10^6)=1.15\cdot 10^6\frac{V}{m}"